# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v2

@@ Abstract
Preconditioned Proximal Point (PPP) algorithms provide a unified framework for splitting methods in image restoration. Recent advancements with RED (Regularization by Denoising) and PnP (Plug-and-Play) priors have achieved state-of-the-art performance in this domain, emphasizing the need for a meaningful particular solution. However, degenerate PPP algorithms typically exhibit weak convergence in infinite-dimensional Hilbert space, leading to uncertain solutions. To address this issue, we propose the Halpern-type Preconditioned Proximal Point (HPPP) algorithm, which leverages the strong convergence properties of Halpern iteration to achieve a particular solution. Based on the implicit regularization defined by gradient RED, we further introduce the Gradient REgularization by Denoising via HPPP called GraRED-HP3 algorithm. The HPPP algorithm is shown to have the regularity converging to a particular solution by a toy example. Additionally, experiments in image deblurring and inpainting validate the effectiveness of GraRED-HP3, showing it surpasses classical methods such as Chambolle-Pock (CP), PPP, RED, and RED-PRO.
@@
Halpern iteration, Preconditioned proximal point algorithms, RED, Image restoration
@@ 1 Introduction
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization (or prior) to mitigate the ill-posedness of IR problems. Examples include total variation (TV) regularization $\left\|\nabla\mathbf{x}\right\|_{1}(\mathbf{K}=\nabla)$ as in [34], and sparsity regularization $\left\|\mathbf{x}\right\|_{1}(\mathbf{K}=I)$.
@@
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [4, Chapter 3]. Following [20, 5], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we can reformulate (2) as:
@@
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [1, 2]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [27]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and has been proven to converge weakly [32]. Since the resolvent $(I+\mathcal{A})^{-1}$ is generally difficult to compute, splitting methods have been proposed to address this issue. The well-known Douglas-Rachford splitting (DRS) [16] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which the resolvents $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another popular splitting algorithm is the Chambolle-Pock (CP) algorithm [10, 30], which solves the saddle-point problem of (1), i.e.,
@@
where $g^{*}:\mathcal{K}\to\mathbb{R}^{+}$ is the Fenchel conjugate functional. He and Yuan [21] were the first to analyze the Primal-Dual Hybrid Gradient (PDHG) by the preconditioned proximal point (PPP) method, given by $\mathbf{0}\in\mathcal{A}\mathbf{u}^{k+1}+\mathcal{M}(\mathbf{u}^{k+1}-\mathbf{u}^{k})$ with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Bredies et al. [6, 7, 5] developed a unified degenerate PPP algorithmic framework with a semi-definite preconditioner. Assuming that $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$ has a full domain and is single-valued, the PPP method with the classic Krasnosel’skii-Mann (KM) iteration solves for the fixed point of $\mathcal{T}$, i.e., $\mathcal{T}(\mathbf{u})=\mathbf{u}$. From the perspective of degenerate PPP [5], both DRS and CP splitting algorithms can be considered as special cases of the PPP algorithm with proper preconditioner.
However, PPP algorithms with a semi-definite preconditioner generally exhibit weak convergence in Hilbert space, resulting in uncertain solutions. It is meaningful to study the limit point of sequences generated by PPP, Bauschke et al. obtained strong convergence results of PPP in the special case of linear relation on $\mathcal{A}$ [2], the fixed-point set $\mathrm{Fix}(\mathcal{T})$ is a linear subspace, the weak limit of the PPP sequence is the $\mathcal{M}$-projection of initial point. The classic Halpern iteration [19] offers the advantage of strong convergence over other iterations (such as the KM iteration) in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [22]. This method is also known as an implicit regularization technique [15, 31]. Due to this implicit bias (projection onto solutions), degenerate PPP algorithms incorporating Halpern iteration are more likely to approximate the true solution, yielding better recovery results for IR problems. Besides the benefit of converging to a particular fixed point, Halpern iteration also possesses acceleration properties widely utilized in machine learning [15, 29]. Even without theoretical guarantees, literature [29] has demonstrated that PDHG with Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Sun et al. propose a globally convergent accelerated degenerate Proximal Point Method (dPPM) by unifying the Halpern iteration and the fast KM iteration. They further accelerate a preconditioned Alternating Direction Method of Multipliers (pADMM) via dPPM [37].
Recently, first-order optimization algorithms (steepest decent (SD), fixed-point (FP) iteration and ADMM) have played an critical role in implicit regularization, such as PnP priors [39] and RED [33], which have achieved state-of-the-art performance in IR problems [8, 44, 25, 38]. PnP methods by replacing proximal operators with denoisers do not seek the minimization of an explicit objective function, which strongly limits their interpretation [14, 25]. The RED prior defines a convex function whose gradient corresponds to the denoising residual itself, formulating a clear and well-defined objective function [14]. Thus, the gradient of RED can be used by these first-order optimization methods. To prove the global convergence of RED, based on fixed-point projection, Cohen et al. proposed the RED-PRO model [14] to bridge between RED and PnP using the hybrid steepest descent method (HSD) [42]. Assuming that the data term $f$ is a strong convexity function and the residual of the denoising neural network is nonexpansive, Ryu et al. proved the convergence of PnP-FBS(Forward-Backward Splitting) and PnP-ADMM by the classical Banach contraction principle [35], but imposing strong convexity on $f$ excludes many IR tasks [25]. To overcome the strong convexity assumption, Hurault et al. proposed proximal denoisers of PnP methods, and the gradient step denoiser [13, 24] can be interpreted as a proximal operator of implicit regularization [25]. However, these works did not obtain the particular solution. It is crucial to obtain a unique, stable, and particular solution for IR problems [17].
Since the Halpern iteration can natually obtain the particular solution (initial point projection onto the fixed-point set). Therefore, Halpern-type preconditioned proximal point called HPPP (5) is proposed to takle the above issue, i.e.,
(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}(\mathbf{u}^{k}),$
where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ is a $\mathcal{M}$-FNE operator (7) and $\{\mu_{k}\}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\mu_{k}\to 0(k\to\infty)$. In addition to TV, implicit regularization such as PnP prior is also suitable for HPPP. Under the nonexpansive assumption, we find that the residual of the denoiser can also be interpreted as the proximal operator of the implicit regularization via Moreau decomposition, making it straightforward to integrate into the primal-dual algorithm with the special semi-definte preconditioner $\mathcal{M}$. These splitting algortihms can be viewed as fixed-point iterations from the perspective of PPP. Based on HPPP and implicit regularization,  [^1] Since the gradient of RED is exactly the residual of the denoiser, we used the GraRED notion to denote the proximal operator of the implicit regularization.we propose the implicit Gradient Regularization by Denoising via HPPP called GraRED-HP3(see Algorithm 1) for IR problems. Not only can it converge to a particular solution, but it also leverages the data adaptivity of implicit regularization. This integration enriches the theoretical and algorithmic understanding of RED and PnP priors. The main contributions are as follows:
@@
    Theoretically, under the semi-definite preconditioner, we analyze the convergence of HPPP in Hilbert space. Compared with PPP, the proposed HPPP has the advantage of converging to a unique particular solution $\mathbf{u}^{*}={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(Lemma A.4), which extends the results of the classic Halpern iteration. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by HPPP; we establish an $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate for $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
@@
    Based on the special preconditioner $\mathcal{M}$, the primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}$. We apply TV or implicit regularization to HPPP, then propose the GraRED-HP3 algorithm for image restoration problems, further discuss the relationship between PnP-ADMM, GraRED-P3 (GraRED via PPP), and the proposed algorithm.
@@
    We numerically verify the regularity of HPPP with a simple 1D example, and further demonstrate the state-of-the-art performance of GraRED-HP3 on image deblurring and inpainting experiments. The results show that GraRED-HP3 outperforms the classic CP [10] and PPP [5] algorithms with TV regularization, as well as the RED [33] and RED-PRO [14] algorithms with TNRD denoiser [12].
The rest of this paper is organized as follows. In section 2, we review some useful preliminaries for convergence analysis, including the background of PPP and implicit gradient RED. In section 3, we establish the convergence and convergence rate of HPPP and apply it to TV regularization. Based on the implicit gradient RED, we propose the GraRED-HP3 algorithm for IR problems. In section 4, we verify the regularity and efficiency of the proposed algorithms with a 1D toy example. Furthermore, we validate the performance of the proposed GraRED-HP3 algorithm through image deblurring and inpainting experiments. Finally, conclusions are presented in section 5.
@@ 2.1 Preconditioned proximal point
Bredies et al. [5] introduced a linear, bounded, self-adjoint and positive semi-definite operator admissible preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates $(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$.
@@
is single-valued, of full-domain, and Lipschitz continuous.
Therefore, the preconditioned proximal point iteration is written into
@@
If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly non-expansive operator [1, section 23] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is semi-definite, the property of $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [6, 9, 41, 5], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
The following notion extends monotone characteristics, i.e., $\mathcal{M}-$monotonicity.
@@
According to [5, Lemma 2.6], if $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
@@
By the KM iteration, the relaxed PPP algorithm is written into
@@
where $\lambda_{k}$ satifies $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 Implicit Gradient RED
Romano et al. introduced an explicit regularizer called Regularization by Denoising (RED) [33] to construct an explicit objective function, which is defined by
(9) $g(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser ($\sigma>0$ denotes a noise level). If $D_{\sigma}(\mathbf{x})$ satisfies homogeneity $D_{\sigma}(c\mathbf{x})=cD_{\sigma}(\mathbf{x})(\forall c>0)$ and has symmetric Jacobian, then $\nabla g(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})$. Under the following nonexpansive Assumption (A[35] on residul $R=I-D_{\sigma}$, i.e.,
(A) $\left\|(\mathbf{x}-D_{\sigma}(\mathbf{x}))-(\mathbf{y}-D_{\sigma}(\mathbf{y}))\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n},$
we will prove that there exists an implicit regularization $\phi$ such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$. Here is the definition of the proximal operator and characterization of proximal operators.
Definition 2.3.
[4, Section 6] Given a function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$, the proximal operator of $\phi(\mathbf{x})$ is defined by

### v4

@@ Abstract
Recently, the degenerate preconditioned proximal point (PPP) method provides a unified and flexible framework for designing and analyzing operator-splitting algorithms such as Douglas-Rachford (DR). However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants. To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method. Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors such as Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method. Finally, numerical experiments including several toy examples and image restoration validate the effectiveness of our proposed algorithms.
@@
Halpern iteration, Preconditioned proximal point algorithms, Plug-and-Play Prior, Regularization by Denoising, Image restoration
@@ 1 Introduction
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization such as TV [48] to mitigate the ill-posedness of IR problems.
@@
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3]. Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as:
@@
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [2, 3]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46]. However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue. The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another way to solve (1) is to transform it into the saddle-point problem [16, 43], i.e.,
@@
where $g^{*}:\mathcal{K}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$. In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) splitting method [16, 43] and the Primal-Dual Hybrid Gradient (PDHG) method [27], are designed and studied extensively. In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Recently, Bredies et al. [11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semi-definite $\mathcal{M}$. By choosing appropriate preconditioners, this framework could cover the DR and CP splitting algorithms [9].
It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34]. For example, Bauschke et al. demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3]. In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point. Moreover, Sun et al. [52] proposed an accelerated preconditioned Alternating Direction Method of Multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6]. Similarly, Chen et al. introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semi-proximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38]. A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8]. Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces. Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29]. Then the Halpern iteration is also known as an implicitly regularized method [22]. Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems. Achieving stable reconstruction for ill-posed inverse problems is important [24]. Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29]. Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Based on the Halpern iteration, the Halpern-type preconditioned proximal point algorithm called HPPP (5) is thus proposed to overcome the two limitations of the degenerate PPP method,
(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k},$
where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53]. Buzzard et al. provided a Consensus Equilibrium interpretation on denoiser priors [14]. Romano et al. introduced RED (Regularization by Denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization. Powerful denoisers (e.g., DnCNN [66]) typically do not meet the conditions of RED [45]. To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45]. Based on the fixed-point projection, Cohen et al. proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model. Meanwhile, Ryu et al. proved the convergence of PnP-FBS and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49]. However, enforcing strong convexity on $f$ precludes many IR tasks [33]. Furthermore, Cohen et al. parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault et al. proposed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33]. Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on their equivalence [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
Based on HPPP and denoiser priors,  [^1] Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.we propose the Gradient Regularization by Denoising via HPPP called GraRED-HP3 for IR problems. The main contributions are as follows:
@@
    Theoretical Contributions. The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semi-definite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution. Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
@@
    Algorithmic Development. We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint. Noting that the PnP-ADMM algorithm can be reformulated as an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
@@
    Experimental Validation. We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency with several toy examples. In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
The rest of this paper is organized as follows. In section 2, we review some preliminaries for convergence analysis. In section 3, we analyze the convergence of HPPP. Combining HPPP and denoiser priors, we propose GraRED-HP3 for IR problems. In section 4, we verify the advantages of HPPP with several toy examples. Furthermore, we validate the performance of GraRED-HP3 through IR experiments. Finally, conclusions are presented in section 5.
@@ 2.1 Preconditioned proximal point
Bredies et al. [9] introduced the degenerate PPP framework with the positive semi-definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
@@
is single-valued, full-domain, and Lipschitz continuous.
Therefore, the PPP iteration is written into
@@
If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly nonexpansive (FNE)operator [2, section 23] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is positive semi-definite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$. The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
@@
The following lemma demonstrates that the firmly non-expansiveness notation can be generalized to the degenerate case [11, 15]. If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
Lemma 2.3 ([9]).
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone. Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
@@
By the KM iteration, the degenerate PPP method is written into
@@
where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 PnP and RED
2.2.1 PnP prior
Venkatakrishnan et al. proposed the first PnP method based on ADMM [56]. PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60]. Sreehari et al. established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51]. When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$. Subsequently, a non-local means denoiser satisfying these conditions was constructed [51].
Algorithm 1 PnP-ADMM
1:Input: Given $\mathbf{u}^{0}=\mathbf{0},\mathbf{x}^{0},\mathbf{z}^{0}$ and total iterations $N>0$
2:for $k=0,1,2,\ldots,N-1$ do
3:     $\mathbf{x}^{k+1}=D_{\sigma}(\mathbf{z}^{k}-\mathbf{u}^{k})$
4:     $\mathbf{z}^{k+1}=\text{prox}_{\lambda f}(\mathbf{x}^{k+1}+\mathbf{u}^{k})$
5:     $\mathbf{u}^{k+1}=\mathbf{u}^{k}+(\mathbf{x}^{k+1}-\mathbf{z}^{k+1})$
6:end for
7:Output: $\mathbf{x}^{N}$
Definition 2.4 ([5]).
Given $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the proximal operator of $\phi(\mathbf{x})$ is defined by